# Bundle Neural Networks for message diffusion on graphs

### v1

@@ 4 Properties of Bundle Neural Networks
In this Section, we give a formal analysis of the BuNN model. In Section 4.1, we derive the fixed points of the bundle heat diffusion, which is the subspace of signals towards which solutions converges exponentially fast, and show that despite converging to fixed points, BuNNs can retain information at the node level and therefore avoid over-smoothing. Section 4.2 discusses how our model can capture long-range interactions and mitigate over-squashing.
@@
A significant limitation of MPNNs is the problem of over-smoothing, the phenomenon of node features becoming increasingly indistinguishable with the increase of the MPNN’s depth. Over-smoothing has been repeatedly highlighted as one of the critical problems in GNNs, making training deep GNNs much more challenging. Such behavior is related to the diffusion used in MPNNs being similar to that of heat diffusion on graphs [23], which converges to uninformative fixed points [^1] Up to degree scaling if using the symmetric-normalized Laplacian, see Li et al. [4]. resulting in a loss of information at the node level. Bodnar et al. [13] show that the richer sheaf (or bundle) structure, however, gives rise to more interesting limiting behavior. Indeed, by Lemma 3.1, since $\lim_{t\rightarrow\infty}\mathcal{H}(t,\ v,\ u)=\frac{d_{u}}{2|\mathsf{E}|}$, the limit over time of a solution is $\frac{1}{2|E|}\sum_{u\in\mathsf{V}}d_{u}\mathbf{O}_{v}^{T}\mathbf{O}_{u}\mathbf{x}_{u}$. To better understand the space of fixed points, notice that any signal $\mathbf{Y}\in\mathbb{R}^{n\times d}$ expressible as $\mathbf{y}_{v}=\frac{1}{2|E|}\sum_{u\in\mathsf{V}}d_{u}\mathbf{O}_{v}^{T}\mathbf{O}_{u}\mathbf{x}_{u}$ for some $\mathbf{X}$ is a fixed point. For any two nodes $u,\ v$ we get:
@@
Consequently, the fixed points of vector diffusion have a global geometric dependency where all nodes relate to each other by some orthogonal transformation, e.g. the output in Figure 1.
Equation 5 provides insight into the smoothing properties of BuNNs. When the bundle is trivial, the equation reduces to $\mathbf{y}_{v}=\mathbf{y}_{u}$, with no information at the node level, i.e. over-smoothing. When it is non-trivial, the output signal can vary across the graph (e.g., $\mathbf{y}_{v}\neq\mathbf{y}_{u}$ for two nodes $u$ and $v$). We summarize this in Proposition 4.1.
@@
While message-passing in MPNNs constitutes a strong inductive bias, this local mechanism is problematic when the task requires the MPNN to capture interactions between distant nodes. These issues have been attributed mainly to the over-squashing of information [7]. Topping et al. [8] and Di Giovanni et al. [9] formalize over-squashing through a sensitivity analysis, giving upper bounds on the partial derivatives of the outputs of the MPNN with respect to the input. In particular, they show that under weak assumptions on the message function, the Jacobian of MPNNs satisfies
$|\partial\left(\mathrm{MPNN}_{\Theta}(\mathbf{X})\right)_{u}/\partial\mathbf{x}_{v}|\leq c^{r}\left(\mathbf{A}^{r}\right)_{uv},$ (6)
for any nodes $u,\ v$, where $r$ is the depth of the network and $c$ is a constant. Given two nodes $u,v$ at a distance $r$, message-passing will require at least $r$ layers for the two nodes to communicate and overcome under-reaching. If $r$ is large, distant nodes can communicate, but Di Giovanni et al. [9] show that over-squashing becomes dominated by vanishing gradients. Built on top of such sensitivity analysis, we compute the Jacobian for BuNN:
@@
Let BuNN be a linear layer defined by Equations 23 with hyperparameter $t$. Then, for any connected graph and nodes $u,\ v$, we have
The form of the Jacobian in Lemma 4.2 differs significantly from the usual form in Equation 6. First, all nodes communicate in a single BuNN layer since $\mathcal{H}(t,u,v)>0$ for all $u,v$, and $t$, allowing for direct pair-wise communication between nodes, making a BuNN layer operate globally similarly to the attention mechanism in the Transformer model, and therefore overcome under-reaching. Secondly, taking $t$ to be large allows BuNNs to operate on a larger scale, allowing stronger communication between distant nodes and overcoming over-squashing without the vanishing gradient problem. Indeed, $\mathcal{H}(t,u,v)$ converges to $\frac{d_{u}}{2|\mathsf{E}|}$ exponentially fast with $t\rightarrow\infty$, showing that for larger scales the sensitivity between two nodes does not depend on how they are connected.
Further, Lemma 4.2 gives a finer picture of the capabilities of BuNNs. For example, to mitigate over-squashing a node may decide to ignore information received from certain nodes while keeping information received from others. This allows the model to reduce the receptive field of certain nodes. We formalize such a behavior in Corollary 4.3.
@@
Consider n nodes $u,\ v,\ \text{and}\ w_{i},\text{ for}\ i=1,\ldots n-2$, of a connected graph with $2$ dimensional bundle such that $\mathbf{x}_{v}\neq\mathbf{x}_{w_{i}}$ and $\mathbf{x}_{u}\neq\mathbf{x}_{w_{i}}\ \forall i$. Then in a BuNN layer, at a given channel, the node $v$ can learn to ignore the information from all $w_{i}s$ while keeping information from $u$.
@@ 5 Expressivity of the model
In this Section, we study the expressive power of BuNNs at the node level from a function approximation perspective. We note that such expressive power is strictly stronger than the more common graph-level expressivity, which usually relates GNNs to the Weisfeiler-Leman hierarchy. In node-level approximation, we desire our model to be able to approximate arbitrary feature transformations over the nodes. Further, existing work generally focuses on a fixed graph or graphs of bounded size [24, 25, 26]. In contrast, we focus on the more challenging case of potentially infinite families of graphs, known as uniform approximation [27]. Such analysis studies the class of feature transformation that GNN architectures can approximate over a potentially infinite collection of graphs $\mathcal{G}$. Given input and output dimensions $c$ and $c^{\prime}$, a family of GNNs will parameterize a subset of continuous feature transformation $\mathcal{C}(\mathcal{G},\mathbb{R}^{c},\mathbb{R}^{c^{\prime}})$ over $\mathcal{G}$. It is natural to ask which feature transformations a specific family of GNNs can or cannot parameterize. To this end, we introduce the notion of compact uniform approximation as an amendment to the definition of uniform approximation from Rosenbluth et al. [27] and discuss our specific choice in the Appendix (Section  D).
@@
As we show in Proposition 5.2, bounded-depth MPNNs fail to have this property for any arbitrary $\mathcal{G}$, since fixing the depth of the MPNNs to $\ell$, we can take a single compactly featured graph $\mathsf{G}\in\mathcal{G}$ with diameter larger than $\ell$. As there is a node whose receptive field does not include all nodes in such a $\mathsf{G}$, the architecture cannot uniformly approximate every function on $\mathcal{G}$ [^2] This phenomenon in MPNNs is often called ‘under-reaching’.
@@
It was shown in Rosenbluth et al. [28] that even with injective Positional Encodings (PEs), there are functions that MPNNs with virtual-nodes cannot approximate uniformly while transformers can, and vice-versa. Instead, we prove that our model is compactly uniformly universal given injective PEs. To the best of our knowledge, this is the first positive uniform approximation result for a GNN, demonstrating the remarkable modeling capabilities of BuNNs.
@@
Let $\mathcal{G}$ be a set of connected graphs with injective positional encodings. Then $2$-layer deep BuNNs with encoder/decoder have compact uniform approximation over $\mathcal{G}$.
@@ 6 Experiments
In this Section, we evaluate our model through a range of synthetic and real-world experiments. We first validate the theory from Section 4 with two synthetic tasks. We then evaluate BuNNs on popular real-world benchmarks. We use truncated Taylor approximations for small values of $t$ ($\leq 1$), while for larger $t$, we use the exact spectral solution. We provide supplementary information on the implementation and precise experimental details in the Appendix (Sections E and F respectively).
@@ 6.1 Synthetic experiments: over-squashing and over-smoothing
Tasks. In this experiment, we compare BuNNs to MPNNs on two new datasets. The tasks are node-regression tasks, in which nodes must average the input features of a subset of nodes as illustrated in Figure 3. The input graph contains two types of nodes, denoted $A$ and $B$, whose features are sampled from disjoint distributions. The target for nodes in $A$ is to output the average input feature over nodes in $B$ and vice-versa.
Figure 3: Synthetic over-squashing (left) and over-smoothing (right) tasks on barbell and clique graphs, respectively. In both cases, the blue nodes output the average over the red nodes and vice-versa.
@@
As a first experiment, to evaluate the capacity to mitigate over-squashing, we consider the case where the underlying graph is a barbell graph consisting of two fully connected graphs - each being a type and bridged by a single edge. This bottleneck makes it hard to transfer information from one cluster to another. As a second experiment, we test the capacity to mitigate over-smoothing by considering the fully connected graph, in which all nodes are connected. The fully connected graph is a worst-case scenario for over-smoothing since after one step of message passing, the features are already fully averaged over the graph. This over-smoothing makes it challenging to predict different outputs over the different clusters.
@@
Setup. As a first baseline, we consider a constant predictor always predicting 00, the expected mean over the whole graph, making it unaware of any cluster-specific information. As a second baseline, we consider the cluster-specific constant predictor predicting the expected mean over the opposite cluster, that is, ±$\frac{\pm\sqrt{3}}{2}$, depending on the cluster. Additionally, we consider GNN baselines to be a node-level MLP, GCN [15], GraphSAGE [29], GAT [30], and NSD [13]. The width of MPNNs is fixed to the minimal depth to avoid under-reaching, namely $3$ for the Barbell and $1$ for the fully connected graph, and ensure the width is large ($>128$) considering the task. We compare these to a BuNN with an MLP learning the bundle maps of a comparable number of parameters and the same depth. We use Adam optimizer with $10^{-3}$ learning rate, batch size $1$, and train for $500$ epochs. We use $100$ samples for training and $100$ samples for testing.
@@
Results. The results for $N=10$ are reported in Table 2. All MPNNs perform poorly on the over-squashing task on the barbell graph. All MPNNs perform comparably to baseline $2$, which is aware of nodes in the same clusters but not the other, showing their incapability to transfer information between clusters. This is explained by the fact that the resistance distance between nodes from different clusters is $1.0$ while it is $0.2$ for nodes in the same cluster, making communication between nodes from different clusters hard. Indeed, resistance distance tightly connects to over-squashing, and the two nodes’ communication strength is inversely proportional to this metric Di Giovanni et al. [9]. On the other hand, BuNN achieves almost perfect accuracy on this task, which supports the claim that BuNN mitigates over-squashing. We note, however, that to solve the task perfectly, we need $t\geq 10$, allowing BuNN to operate on a larger scale more adapted to the task. To solve this task, BuNN can assign the orthogonal maps to separate the two types of nodes, making each node listen only to the nodes of the other type, a behavior proved to be possible in Corollary 4.3.
Task
@@
Table 2: Mean squared error (MSE) of different models for the over-squashing and over-smoothing experiments on the barbell and clique graphs, respectively.
Similarly, the over-smoothing task on the clique graph is also challenging. Indeed, GCN and GAT perform exceptionally poorly, comparably to Baseline $1$, which only has access to global information and not cluster-specific information. To the best of our knowledge, these are the only models with formal proofs of over-smoothing [6, 31] respectively. GraphSAGE performs slightly better because it processes neighbors differently than it processes nodes themselves, but it is still close to node-level MLPs and Baseline 2. Moreover, NSD solves the task as expected due to their capability to mitigate over-smoothing. Lastly, BuNN also learns to solve this task due to its capacity to mitigate over-smoothing. Indeed, it can learn to ignore nodes from a given cluster, a behavior proved to be possible in Corollary 4.3.

### v2

@@ 4 Properties of Bundle Neural Networks
We now give a formal analysis of the BuNN model. In Section 4.1, we derive the fixed points of the bundle heat diffusion, which the subspace of signals towards which solutions converges, and show that even in the limiting case BuNNs can retain information at the node level and therefore avoid over-smoothing. Section 4.2 discusses how our model can capture long-range interactions and mitigate over-squashing.
@@
A major limitation of MPNNs is over-smoothing, where node features become indistinguishable as MPNN depth increases. This phenomenon is a major challenge for training deep GNNs. It arises because the diffusion in MPNNs resembles heat diffusion on graphs (Di Giovanni et al., ), which converges to uninformative fixed points [^1] Up to degree scaling if using the symmetric-normalized Laplacian, see Li et al. (2018)., leading to a loss of information at the node level. Bodnar et al. (2022) show that the richer bundle structure, however, gives rise to richer limiting behavior. Indeed, by Lemma 3.1, since $\lim_{t\rightarrow\infty}\mathcal{H}(t,\ v,\ u)=\frac{d_{u}}{2|\mathsf{E}|}$, the limit over time of a solution is $\frac{1}{2|E|}\sum_{u\in\mathsf{V}}d_{u}\mathbf{O}_{v}^{T}\mathbf{O}_{u}\mathbf{x}_{u}$. To understand the space of fixed points, notice that any $\mathbf{Y}\in\mathbb{R}^{n\times d}$ expressible as $\mathbf{y}_{v}=\frac{1}{2|E|}\sum_{u\in\mathsf{V}}d_{u}\mathbf{O}_{v}^{T}\mathbf{O}_{u}\mathbf{x}_{u}$ for some $\mathbf{X}$ is a fixed point, and for any two nodes $u,\ v$:
@@
Consequently, the fixed points of vector diffusion have a global geometric dependency where all nodes relate to each other by some orthogonal transformation, e.g. the output in Figure 3. Equation 5 provides insight into the smoothing properties of BuNNs. When the bundle is trivial, the equation reduces to $\mathbf{y}_{v}=\mathbf{y}_{u}$, with no information at the node level, i.e. over-smoothing. When it is non-trivial, the output signal can vary across the graph (e.g., $\mathbf{y}_{v}\neq\mathbf{y}_{u}$ for two nodes $u$ and $v$). We summarize this in Proposition 4.1 and prove an implication in terms of over-smoothing in Proposition 4.2. Similar results showing that the signal can survive in deep layers as opposed to resulting in the constant signal have previously been done empirically and theoretically for other architectures, e.g. Chamberlain et al. (2021); Fu et al. (2023).
@@
Proposition 4.2.
Let $\mathbf{X}$ be $2$-dimensional features on a connected $\mathsf{G}$, and assume that there are two nodes $u,\ v$ with $\mathbf{p}_{u}\neq\mathbf{p}_{v}$. Then there is an $\delta>0$ such that for every $K$, there exist a $K$-layer deep BuNN with $2$ dimensional stalks, $t=\infty$, ReLU activation, $\phi^{(\ell)}$ is a node-level MLP, and $\|\mathbf{W}^{(\ell)}\|\leq 1$, with output $\mathbf{Y}$ such that $\|\mathbf{y}_{v}-\mathbf{y}_{u}\|>\delta$.
@@
While message-passing in MPNNs constitutes a strong inductive bias, it is problematic when the task requires the MPNN to capture interactions between distant nodes. These issues have been attributed mainly to the over-squashing problem. Topping et al. (2022) and Di Giovanni et al. (2023) formalize over-squashing through a sensitivity analysis, giving upper bounds on the sensitivity of the output at a node with respect to the input at an other. In particular, they show that under weak assumptions on the message function, the Jacobian of an $\mathrm{MPNN}$ satisfies the following inequality
$|\partial\left(\mathrm{MPNN}_{\Theta}(\mathbf{X})\right)_{u}/\partial\mathbf{x}_{v}|\leq c^{\ell}\left(\mathbf{A}^{\ell}\right)_{uv},$ (6)
for any nodes $u,\ v$, where $\ell$ is the depth of the network and $c$ is a constant. Given two nodes $u,v$ at a distance $r$, message-passing will require at least $r$ layers for the two nodes to communicate and overcome under-reaching. If $\ell$ is large, distant nodes can communicate, but Di Giovanni et al. (2023) show that over-squashing becomes dominated by vanishing gradients. Building on top of such sensitivity analysis, we compute the Jacobian for a BuNN layer in Lemma 4.3.
@@
Let BuNN be a linear layer defined by Equations 1, 23 with hyperparameter $t$. Then, for any connected graph and nodes $u,\ v$, we have
$\frac{\partial\left(\operatorname{BuNN}\left(\mathbf{X}\right)\right)_{u}}{\partial\mathbf{x}_{v}}=\mathcal{H}(t,u,v)\mathbf{O}_{u}^{T}\mathbf{W}\mathbf{O}_{v}.$
The form of the Jacobian in Lemma 4.3 differs significantly from the usual form in Equation 6. First, all nodes communicate in a single BuNN layer since $\mathcal{H}(t,u,v)>0$ for all $u,v$, and $t$, allowing for direct pair-wise communication between nodes, making a BuNN layer operate globally similarly to Transformer model, and therefore overcome under-reaching. Secondly, taking $t$ to be large allows BuNNs to operate on a larger scale, allowing stronger communication between distant nodes and overcoming over-squashing without the vanishing gradient problem.
Further, Lemma 4.3 gives a finer picture of the capabilities of BuNNs. For example, to mitigate over-squashing a node may decide to ignore information received from certain nodes while keeping information received from others. This allows the model to reduce the receptive field of certain nodes:
@@
Consider n nodes $u,\ v,\ \text{and}\ w_{i},\text{ for}\ i=1,\ldots n-2$, of a connected graph with $2$ dimensional bundle such that $\mathbf{p}_{v}\neq\mathbf{p}_{w_{i}}$ and $\mathbf{p}_{u}\neq\mathbf{p}_{w_{i}}\ \forall i$. Then in a BuNN layer with MLP $\phi$, at a given channel, the node $v$ can learn to ignore the information from all $w_{i}s$ while keeping information from $u$.
@@ 5 Expressivity of the model
We now characterize the expressive power of BuNNs from a feature transformation perspective. This analysis extends that of Bodnar et al. (2022), who show that heat diffusion on sheaves can be expressive enough to linearly separate nodes in the infinite time limit. Instead, our results concern the more challenging problem of parameterizing arbitrary feature transformations and hold for finite time.
Most work on GNN expressivity characterize the ability of GNNs to distinguish isomorphism classes of graphs or nodes (Xu et al., 2019b; Morris et al., 2019; Azizian & Lelarge, 2021; Geerts & Reutter, 2022) or equivalently to approximate functions on them (Chen et al., 2019). These results typically rely on two major assumptions: 1) node features are fixed for each graph or are ignored completely; and 2) apply to a single graph or finitely many graphs of bounded size. Instead, our setting 1) includes features ranging in an infinite uncountable domain, and 2) includes infinite families of graphs. To this end, we define the notion of compact uniform approximation as a modification to that of uniform approximation from Rosenbluth et al. (2023) and we discuss our choice in Appendix C.
@@
Injective Positional Encodings (PEs). In the case of a finite collection of graphs $\mathcal{G}$ and a fixed finite feature space $K$, the above definition reduces to the setting of Theorem 2 in Morris et al. (2019), since the node features are fixed to the finitely many values in $K$. However, when $K$ is not finite, the arguments in Morris et al. (2019) no longer work, as detailed in Appendix D. Consequently, Definition 5.1 subsumes graph-isomorphism testing, and it is, therefore, too strong for a polynomial-time GNN to satisfy. Hence, we will assume that the GNN has access to injective positional encodings, which we formally define in Definition D.1. This allows us to bypass the graph-isomorphism problem, while not trivializing the problem as we show next.
Negative results with injective PE. Characterizing the expressive power of GNNs in the uniform setting is an active area of research, with mostly negative results. Rosenbluth et al. (2024) prove negative results in the non-compact setting for both MPNNs with virtual-node and graph transformers, even with injective positional encodings. While extending their result to the compact setting is outside the scope of our work, we prove in Proposition 5.2 a negative result for bounded-depth MPNNs with injective PEs. Indeed, fixing the depth of the MPNNs to $\ell$, we can take a single compactly featured graph $\mathsf{G}\in\mathcal{G}$ with a diameter larger than $\ell$. As there is a node whose receptive field does not include all nodes in such a $\mathsf{G}$, the architecture cannot uniformly approximate every function on $\mathcal{G}$ [^2] This phenomenon in MPNNs is often called ‘under-reaching’.
@@
Universality of BuNN. In contrast to the negative results above, we now show that when equipped with injective PEs, BuNNs are universal with respect to compact uniform approximation. To the best of our knowledge, Theorem 5.3 is the first such positive feature approximation result for a GNN architecture, which demonstrates the remarkable modelling capabilities of BuNNs.
@@
Let $\mathcal{G}$ be a possibly infinite set of connected graphs equipped with injective positional encodings. Then $2$-layer deep BuNNs with encoder/decoder at each layer and $\phi^{(1)}$, $\phi^{(2)}$ being $2$-layer deep MLP have compact uniform approximation over $\mathcal{G}$.
@@ 6 Experiments
In this section, we evaluate BuNNs through a range of synthetic and real-world experiments. We first validate the theory from Section 4 with two synthetic tasks. We then evaluate BuNNs on popular real-world benchmarks. We use truncated Taylor approximations for small values of $t$, while for larger $t$, we use the truncated spectral solution. We provide supplementary information on the implementation and precise experimental details in the Appendix (Sections E and F respectively). [^3] All code can be found at https://github.com/jacobbamberger/BuNN
@@ 6.1 Synthetic experiments: over-squashing and over-smoothing
Tasks. In this experiment, we propose two new node-regression tasks in which nodes must average the input features of a subset of nodes. The input graph contains two types of nodes, whose features are sampled from disjoint distributions. The target for nodes is to output the average input feature over nodes of the other type and vice-versa, as illustrated in Figure 4. First, we test the capacity to mitigate over-squashing. We consider the case where the underlying graph is a barbell graph consisting of two fully connected graphs - each being a type and bridged by a single edge. This bottleneck makes it hard to transfer information from one cluster to another. Second, we test the capacity to mitigate over-smoothing. We consider the fully connected graph, in which all nodes are connected. The fully connected graph is a worst-case scenario for over-smoothing since after one step of message passing, the features are fully averaged over the graph and hence over-smoothed.
Setup. As a first baseline, we consider a constant predictor always predicting 00, the expected mean over the whole graph. As a second baseline, the cluster-specific constant predictor predicting the expected mean over the opposite cluster, that is, ±$\frac{\pm\sqrt{3}}{2}$ depending on the cluster.
Figure 4: Synthetic over-squashing (left) and over-smoothing (right). In both cases, blue nodes output the average over the red nodes and vice-versa.
@@
Additionally, we consider GNN baselines to be a node-level MLP, GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), NSD (Bodnar et al., 2022), and a fully connected GraphGPS (Rampášek et al. (2022)). The depth of MPNNs is fixed to the minimal depth to avoid under-reaching, namely $3$ for the barbell and $1$ for the fully connected graph, and ensure the width is large ($>128$) considering the task. The depth of the transformer is tuned between $1$ and $3$ with $2$ heads. We compare these to a BuNN with an MLP learning the bundle maps of a comparable number of parameters and the same depth. We use Adam optimizer with $10^{-3}$ learning rate, batch size $1$, and train for $500$ epochs. We use $100$ samples for both training and testing.
@@
Results. The results for $N=10$ are reported in Table 2. All MPNNs perform poorly on the over-squashing task. All MPNNs perform comparably to baseline $2$, showing their incapability to transfer information between clusters. This is explained by the fact that nodes from different clusters have high commute time and effective resistance, which tightly connects to over-squashing Di Giovanni et al. (2023); Black et al. (2023); Dong et al. (2024). On the other hand, BuNN achieves almost perfect accuracy on this task, which supports the claim that BuNN mitigates over-squashing. We note, however, that to solve the task perfectly, we need $t\geq 10$, allowing BuNN to operate on a larger scale more adapted to the task. To solve this task, BuNN can assign the orthogonal maps to separate the two types of nodes, making each node listen only to the nodes of the other type, a behavior proved to be possible in Corollary 4.4.
Table 2: BuNN mitigates over-smoothing and over-squashing. Mean squared error (MSE) of different models on the two synthetic tasks.
@@
Similarly, the over-smoothing task on the clique graph is also challenging. Indeed, GCN and GAT perform exceptionally poorly, comparably to Baseline $1$ which only has access to global information. Indeed, to the best of our knowledge, these are the only models with formal proofs of over-smoothing (Cai & Wang, 2020; Wu et al., 2023). GraphSAGE performs slightly better because it processes neighbors differently than it processes nodes themselves. Interestingly, GPS fails to generalize providing evidence for over-smoothing and over-squashing in graph transformers. Moreover, NSD and BuNN solve the task due to their capability to mitigate over-smoothing. Indeed, BuNN can learn to ignore nodes from a given cluster as proved in Corollary 4.4.